How many molecules of NH4Br are in 3.406g of NH4Br?

1
Expert's answer
2021-10-31T12:42:52-0400

Molar mass of NH4Br= 97.94

3.406/97.94= 0.0348mol

=0.0348×6.023×1023

=2.09×1022molecules